# Causality-enhanced Decision-Making for Autonomous Mobile Robots in Dynamic Environments

### v3

@@ 3.3 Inference Pipeline
Once built, the causal inference engine can be used by the robot to make decisions based on inferred information. Specifically, the inference pipeline begins with a query from the robot at the start of an assigned task. Before proceeding, the robot must first gather information that is crucial for successfully accomplishing the task. As previously mentioned, this information can be related to the environment in which the robot operates or to its internal state, both of which influence task completion.
The robot’s query is formulated as an intervention, representing an action it considers taking to accomplish the task (e.g., What if I do(X)?). Through the causal inference engine, which has been previously learnt, the inference pipeline evaluates the potential consequences of this action by estimating how relevant factors would change as a result of an intervention or condition. The output of this process is the conditional probability distribution of the queried factor, taking into account causal links within the causal model. Once again, this inference step is performed using pyAgrum5, which provides a full implementation of do-calculus [48] for this step [49]. The estimated value of the factor is then obtained by computing its expectation over the inferred probability distribution. The estimated value is finally embedded in a heuristic used by the A* algorithm to define the robot’s path. If a valid path is found under certain conditions, the inference pipeline returns a decision to proceed with the task using the selected path. In contrast, if no valid path is found, an abort signal is returned to the robot. These steps constitute the inference pipeline and are shown in Fig. 2. A use-case implementation of the framework is presented in the following section, including an example of causal inference for decision making in Sec. 4.3.
@@ 4 Simulating Contextual Human-Robot Spatial Interactions
PeopleFlow enables the creation of realistic, diverse, and repeatable spatial interaction scenarios by incorporating contextual factors that dynamically influence human and robot behaviours. These contextual factors—such as time, location, and environmental obstacles—are crucial in shaping the interactions dynamic. In addition, as highlighted in previous work [26, 50], such simulation environments not only support data generation but also provide a controlled setting for verifying the correctness of causal models and evaluating the experimental procedures used to extract them. This helps reduce both time wasted on ineffective experiments and unnecessary financial costs.
@@ 4.1 PeopleFlow
We developed PeopleFlow by building upon our previous work4[26]. The latter is a Gazebo-based simulator to support causal inference in HRSI scenarios, enabling the collection of observational data and allowing for interventions involving both the robot and humans. PeopleFlow is designed for modelling human-human and human-robot spatial interactions in shared workspaces. Its key novelty lies in contextualising agent behaviours—both for humans and robots—based on scenario-specific factors such as time-dependent routines, robot states, and dynamic environmental conditions, all of which influence interaction patterns. The simulator is built on Gazebo and ROS and features a TIAGo [^7] https://pal-robotics.com/robot/tiago/ robot alongside multiple pedestrian agents. Human behaviour is modelled using the pedsim_ros [^8] https://github.com/srl-freiburg/pedsim_ros library, which simulates group and individual movements based on a social force model [51]. On top of this simulation stack, we introduce two new ROS nodes: the Context Manager and the Context-Pedsim Bridge.
Figure 4: (Top) Warehouse-like scenario map with obstacles depicted as black-band rectlangles, and robot target stations as diamonds. Its implementation in our simulator: (Centre) Gazebo view, (Bottom) RViz view.
The Context Manager node governs the scenario-specific contextual factors (e.g. time-dependent routines) and communicates the current context to the Context-Pedsim bridge node. The Context-Pedsim Bridge handles the generation of new tasks for pedestrian agents and computes their navigation plans accordingly. In more details, we introduced a waypoint-based division of the environment into meaningful regions, such as workspaces, canteens, offices, and corridors. These waypoints serve both as semantic labels and as navigation nodes for guiding agent movement. Their connectivity is defined based on their spatial arrangement while ensuring that direct paths between them do not intersect obstacles or fall within the influence radius of other waypoints. When an agent requests a new task, the Context-Pedsim Bridge node, based on the current context, selects a destination waypoint (i.e., an activity station) and assigns a random activity duration. The agent’s current position and assigned destination are then passed to an A*-based path planner, which computes the shortest waypoint sequence to reach the goal. Once the agent reaches the destination, it remains stationary for the designated time to simulate engagement in the activity. Afterward, the task cycle restarts. The whole strategy is shown in Fig. 3. Mapping destinations to specific waypoints and driving task assignment based on contextual rules (e.g., time of day), enables the production of more realistic and interpretable human activities.
@@
4.2 Context-sensitive Humans and Robot Behaviours in a Warehouse Setting
In this paper, we focus on a warehouse-like scenario—based on our funding project DARKO [^9] https://darko-project.eu/ and common industrial settings [52]—as a case study to evaluate our causality-based robot decision-making framework. Fig. 4 (top) illustrates the map of the warehouse environment implemented in our simulator. The environment consists of several main areas, including the entrance, two offices, two toilets, a canteen, a working area, and a corridor connecting all these spaces. Black-band rectangles indicate shelves in the working area where humans perform tasks and tables in the canteen, while diamonds represent target points for the robot (e.g., trays and conveyors). In Gazebo (Fig. 4, centre), humans are represented as red cylinders, whereas in RViz (Fig. 4, bottom), they are visualised as manikins. We now describe the contextual factors, robot features, and how we modelled worker activities within this environment.
@@ 4.2.1 Contextual Factors
  • Time slot ($S$), which affects human activities and robot task execution, simulating different operational dynamics throughout a typical warehouse shift;
@@
  • Presence of static obstacle ($O$), which models the presence of obstacles not included in the map of the environment (e.g., a forklift or a box left by a human). This factor affects robot velocity and battery consumption, as they require additional maneuvers.
@@ 4.2.2 Robot Features
Autonomy is a key factor for any mobile robot operating in real-world environments. Among the many aspects that influence autonomy, battery life is crucial, as it directly affects the robot’s ability to perform tasks and activities. In our simulator, we model battery dynamics, including both charging and discharging processes, to better approximate real-world conditions. Additionally, we model how the presence of an unexpected static obstacle in the environment affects the robot’s battery dynamics and velocity. This allows us to more realistically simulate the robot’s operation in a warehouse environment, where battery consumption is a critical factor that must be considered in the decision-making process [53, 54, 55, 56].
We model the battery level’s change ($L$) based on two contextual factors: the charging status ($C$) and the presence of an unexpected static obstacle ($O$). Inspired by previous works [57], the battery dynamic in discharging mode ($C=0$) is based on two primary factors: static and dynamic consumption. Static battery consumption accounts for energy usage when the robot is idle, primarily due to active sensors and other essential components. In contrast, dynamic battery consumption represents the energy expended due to the robot’s movement. However, battery consumption increases in situations where the robot is forced to stop suddenly or perform maneuvers to avoid an unexpected static obstacle in the environment. Instead, when the robot is in charging mode, the battery level increases at a predefined charging rate. To capture these factors, we model the change in battery level over time using the following equation:
@@
where $L_{t}$ represents the battery level’s change at time step $t$, $\Delta_{t}$ is the time step duration, $K_{s}$ is a constant representing static battery consumption, $K_{d}$ is a coefficient related to battery consumption due to motion, $V_{t}$ is the robot’s velocity at time step $t$, $K_{o}$ is a coefficient representing the increase in battery consumption caused by static obstacles that force the robot to modify its behaviour unexpectedly. $K_{c}$ is the charging rate when the robot is in charging mode. The two contextual factors, $C$ and $O$, indicate the robot’s charging status and the presence of a static obstacle, respectively. $C$ is 00 when the robot is discharging and $1$ when it is charging, while $O$ is 00 when no obstacle is present and $1$ when an obstacle is detected.
@@
where $B_{t}$ and $B_{t-1}$ denote the battery percentage at time $t$ and $t-1$ respectively, and $L_{t-1}$ represents the battery level’s change, as modelled by (1), at time $t-1$.
@@
  • $L\xleftarrow{}O\xrightarrow{}V$: $L$ is affected by $O$. Moreover, $O$ influences $V$ because the robot is forced to stop or maintain a low velocity when encountering an unexpected obstacle.
@@
Although we use time-series data, we omit explicit time indices $t$ and $t-1$ for simplicity, as all the described links represent contemporaneous relationships occurring at time $t$. The considered relationships are illustrated by the Directed Acyclic Graph (DAG) in Fig. 5.
@@ 4.2.3 Worker Activities
Figure 6: Two examples of people congestion areas at different time slots. (Top left) Probability distribution of waypoint choices for people in time slot $2$ (09:00–10:00). (Top right) RViz view showing people primarily concentrated at the waypoints with the highest probabilities from the distribution graph. (Bottom left) Probability distribution of waypoint choices for people during lunchtime ($S6$, 13:00–14:00). (Bottom right) RViz view showing the majority of workers gathered in the canteen area.
In a warehouse, workers are assigned specific tasks at different times of the day. The activities assigned to each person are generally the same every day [58]. For instance, warehouse operators may be responsible for unloading goods in the morning, sorting and shelving items during the day, and preparing shipments in the evening. These tasks are generally repeated every working day. As a result, certain areas of the warehouse experience higher worker traffic during specific time periods, creating congestion zones.
In order to model this phenomenon, we divided the environment into distinct areas (waypoints) and introduced time as a contextual factor to differentiate workers’ activities at different times of the day. Each waypoint represents a specific area of the warehouse that experiences varying levels of congestion depending on the time, making it an essential contextual factor in our model. To capture the temporal dynamics of workers activity, we divided the working period into hourly time slots from 08:00 to 18:00, along with an additional off period. Within each time slot, worker activities are predefined and consistent, reflecting the structured nature of warehouse operations. The 08:00–09:00 slot models workers’ arrival and the beginning of the working day. To enhance realism, we model each worker arriving at a random time within this slot before starting their assigned activity. Time slots from 09:00 to 13:00 represent the normal morning working period, during which workers are typically present in operational areas such as shelves and offices. The 13:00–14:00 slot represents the lunch break, where workers are primarily in the canteen area. Following this, time slots from 14:00 to 18:00 model the afternoon working period, with the 17:00–18:00 slot representing the end of the working day, when workers finish their tasks and leave the warehouse after completing their standard eight-hour shift. After 18:00, we model the off period, during which no workers are present in the warehouse.
An example of a congestion zone in two different time slots is illustrated in Fig. 6. Based on this, we defined a variable called people density ($D$) to quantify the number of people in a specific waypoint. For each waypoint $W$ in the set of waypoints $\Omega$, the density $D$ is defined as follows:
@@
By incorporating both waypoints and time slots as contextual factors, our model effectively represents fluctuating congestion levels within the warehouse. This enables a more accurate simulation of human-robot interactions, where robot navigation decisions account for dynamically changing environmental conditions. Following this, the causal influence from the contextual factors time slot $S$ and waypoint $W$ on the people density is $W\xrightarrow{}D\xleftarrow{}S$. Again, for simplicity, we have omitted the explicit time indices $t$ and $t-1$. The latter causal relations are added to the model in Fig. 5.
@@ 4.3 Causality-based Robot Decision Making in PeopleFlow
Our causality-based decision making framework presented in Sec. 3.3 is a general causal inference approach adaptable to various scenarios. Here, we give more details on the factors that we want to infer with it and how they are used in the A* algorithm to decide whether to execute or not a given task, and in case to determine the best path.
@@
Given the robot and workers features extracted from the simulator, as well as the context factors explained in Sec. 4.2.1, the crucial aspects affecting robot task completion are the robot battery consumption $L$ and the people density $D$ in specific areas of the environment. Given their dependencies with other variables and/or contextual factors, the robot can infer $L$ based on its velocity $V$ and charging status $C$, while $D$ can be estimated using the specific time slot $S$ and waypoint $W$. These inferences are made under the assumption that the robot has knowledge of its own velocity and charging status. Additionally, in a warehouse shared with workers, we assume that the robot is aware of the current time, enabling it to determine $S$, as well as the predefined waypoints in the environment.
For instance, consider a scenario where the robot is assigned the task “transport an object to the delivery point”, as depicted in Fig. 2. To determine in advance whether the task can be successfully completed or if it should be immediately refused, the robot poses the following query:
“What if I go to the delivery point now at velocity $v$?”
@@
To clarify, the robot queries the causal inference engine to determine whether it can start the task at the current $S$ and travel at velocity $v$. Hence, for each waypoint in $\Omega$, these two interventions allow the causal inference engine to estimate the expected values of $L$ and $D$, which represent the cost associated with the specific waypoint in terms of battery consumption and the risk of encountering people or getting stuck, respectively. These quantities are then used as cost components for navigation planning. In particular, the battery cost for traversing an arc between two waypoints is first computed by determining the traversal time, which is the distance between the waypoints divided by the velocity $v$ (used in the intervention). This duration is then multiplied by the expected $\hat{L}$ from Eq. (6), which corresponds to the battery consumption per time step when moving at velocity $v$. The result is the expected battery cost for that arc. These estimated values, along with the distance between waypoints, are finally used in a heuristic function defined below.
@@
$h(w_{i})=\sum\limits_{i=1}^{n-1}\Big{(}\lambda_{D}\cdot\delta(w_{i},w_{i+1})+\lambda_{D}\cdot\hat{D}(w_{i})+\lambda_{L}\cdot|\hat{L}(w_{i},w_{i+1})|\Big{)}$ (9)
where $\delta(w_{i},w_{i+1})$ represents the Euclidean distance between two consecutive waypoints, $\hat{D}(w_{i})$ is the expected people density at waypoint $w_{i}$ from the causal inference engine, $\hat{L}(w_{i},w_{i+1})$ is the expected battery consumption to reach waypoint $w_{i+1}$ from $w_{i}$, and $\lambda_{D},\lambda_{D},\lambda_{L}$ are weighting coefficients that balance the trade-off between distance minimisation, congestion avoidance, and battery efficiency.

### v4

@@ 3.3 Inference Pipeline
Once built, the causal inference engine can be used by the robot to make decisions based on inferred information. It is important to note that the causal discovery and learning parameters steps are performed only once and are not part of the real-time inference pipeline. Indeed, in this step the robot uses the already-learned causal inference engine to make causality-aware decisions based on new observations, leveraging the relationships encoded in the causal model. Specifically, the inference pipeline begins with a query from the robot at the start of an assigned task. Before proceeding, the robot must first gather information that is crucial for successfully accomplishing the task. As previously mentioned, this information can be related to the environment in which the robot operates or to its internal state, both of which influence task completion.
The robot’s query is formulated as an intervention, representing an action it considers taking to accomplish the task (e.g., What if I do(X)?). Through the causal inference engine, which has been previously learnt, the inference pipeline evaluates the potential consequences of this action by estimating how relevant factors would change as a result of an intervention or condition. The output of this process is the conditional probability distribution of the queried factor, taking into account causal links within the causal model. Once again, this inference step is performed using pyAgrum5, which provides a full implementation of do-calculus (Pearl, 2009) for this step (Ducamp et al., 2020). The estimated value of the factor is then obtained by computing its expectation over the inferred probability distribution. The estimated value is finally embedded in a heuristic used by the A* algorithm to define the robot’s path. If a valid path is found under certain conditions, the inference pipeline returns a decision to proceed with the task using the selected path. In contrast, if no valid path is found, an abort signal is returned to the robot. These steps constitute the inference pipeline and are shown in Fig. 2. A use-case implementation of the framework is presented in the following section, including an example of causal inference for decision making in Sec. 4.3.
@@ 4 Simulating Contextual Human-Robot Spatial Interactions
PeopleFlow enables the creation of realistic, diverse, and repeatable spatial interaction scenarios by incorporating contextual factors that dynamically influence human and robot behaviours. These contextual factors—such as time, location, and environmental obstacles—are crucial in shaping the interactions dynamic. In addition, as highlighted in previous work (Castri et al., 2024a, b), such simulation environments not only support data generation but also provide a controlled setting for verifying the correctness of causal models and evaluating the experimental procedures used to extract them. This helps reduce both time wasted on ineffective experiments and unnecessary financial costs.
@@ 4.1 PeopleFlow
We developed PeopleFlow by building upon our previous work4(Castri et al., 2024a). The latter is a Gazebo-based simulator to support causal inference in HRSI scenarios, enabling the collection of observational data and allowing for interventions involving both the robot and humans. The main limitation of our previous work was the inability to contextualise agent spatial behaviours, preventing the generation of stochastic and realistic human trajectories. Indeed, in ROSCausal_HRISim, the agent’s goal sequence is fully deterministic. PeopleFlow represents its natural extension, and is designed for modelling human-human and human-robot spatial interactions in shared workspaces. Its key novelty lies in contextualising agent behaviours—both for humans and robots—based on scenario-specific factors such as time-dependent routines, robot states, and dynamic environmental conditions, all of which influence interaction patterns. The simulator is built on Gazebo and ROS and features a TIAGo [^7] https://pal-robotics.com/robot/tiago/ robot alongside multiple pedestrian agents. Human behaviour is modelled using the pedsim_ros [^8] https://github.com/srl-freiburg/pedsim_ros library, which simulates group and individual movements based on a social force model (Helbing and Molnar, 1995). On top of this simulation stack, we introduce two new ROS nodes: the Context Manager and the Context-Pedsim Bridge.
The Context Manager node governs the scenario-specific contextual factors (e.g. time-dependent routines) and communicates the current context to the Context-Pedsim bridge node. The Context-Pedsim Bridge handles the generation of new tasks for pedestrian agents and computes their navigation plans accordingly. In more detail, we introduced a waypoint-based division of the environment into meaningful regions, such as workspaces, canteens, offices, and corridors. These waypoints serve both as semantic labels and as navigation nodes for guiding agent movement. Their connectivity is defined based on their spatial arrangement while ensuring that direct paths between them do not intersect obstacles or fall within the influence radius of other waypoints. When an agent requests a new task, the Context-Pedsim Bridge node, based on the current context, selects a destination waypoint (i.e., an activity station) and assigns a random activity duration. The agent’s current position and assigned destination are then passed to an A*-based path planner, which computes the shortest waypoint sequence to reach the goal. Once the agent reaches the destination, it remains stationary for the designated time to simulate engagement in the activity. Afterward, the task cycle restarts. The whole strategy is shown in Fig. 3. Mapping destinations to specific waypoints and driving task assignment based on contextual rules (e.g., time of day), enables the production of more realistic and interpretable human activities.
@@
Figure 4: (a) An illustrative example of an Amazon warehouse setting. (b) Warehouse-like scenario map with obstacles depicted as black-band rectangles, and robot target stations as diamonds. Its implementation in our simulator: (c) Gazebo view, (d) RViz view.
4.2 Context-driven Human-Robot Behaviours in Warehousing
In this paper, we focus on a warehouse-like scenario–inspired by our funding project DARKO [^9] https://darko-project.eu/ and common industrial settings (Dukic and Opetuk, 2012)–as a case study to evaluate our causality-based robot decision-making framework. Fig. 4 illustrates our approach: a typical Amazon warehouse scenario [^10] https://khak.com/amazon-facility-davenport (Fig. 4(a)) inspired the design of our simulated environment map (Fig. 4(b)). The environment consists of several main areas, including the entrance, two offices, two toilets, a canteen, a working area, and a corridor connecting all these spaces. Black-band rectangles indicate shelves in the working area where humans perform tasks and tables in the canteen, while diamonds represent target points for the robot (e.g., trays and conveyors). In Gazebo (Fig. 4(c)), humans are represented as red cylinders, whereas in RViz (Fig. 4(d)), they are visualised as manikins. We now describe the contextual factors, robot features, and how we modelled worker activities within this environment.
@@ 4.2.1 Contextual Factors
  • Time-slot ($S$), which affects human activities and robot task execution, simulating different operational dynamics throughout a typical warehouse shift;
@@
  • Presence of static obstacle ($O$), which models the presence of obstacles not included in the map of the environment (e.g., a forklift or a box left by a human). This factor affects robot velocity and battery consumption, as they require additional manoeuvrers.
@@ 4.2.2 Robot Features
Autonomy is a key factor for any mobile robot operating in real-world environments. Among the many aspects that influence autonomy, battery life is crucial, as it directly affects the robot’s ability to perform tasks and activities. In our simulator, we model battery dynamics, including both charging and discharging processes, to better approximate real-world conditions. Additionally, we model how the presence of an unexpected static obstacle in the environment affects the robot’s battery dynamics and velocity. This allows us to more realistically simulate the robot’s operation in a warehouse environment, where battery consumption is a critical factor that must be considered in the decision-making process (Tomy et al., 2019, 2020; Fragapane et al., 2021; Tosello et al., 2024).
We model the battery level’s change ($L$) based on two contextual factors: the charging status ($C$) and the presence of an unexpected static obstacle ($O$). Inspired by previous works (Hou et al., 2018), the battery dynamic in discharging mode ($C=0$) is based on two primary factors: static and dynamic consumption. Static battery consumption accounts for energy usage when the robot is idle, primarily due to active sensors and other essential components. In contrast, dynamic battery consumption represents the energy expended due to the robot’s movement. However, battery consumption increases in situations where the robot is forced to stop suddenly or perform manoeuvrers to avoid an unexpected static obstacle in the environment. Instead, when the robot is in charging mode, the battery level increases at a predefined charging rate. To capture these factors, we model the change in battery level over time using the following equation:
@@
where $L_{t}$ represents the battery level’s change at time step $t$, $\Delta_{t}$ is the time step duration, $K_{s}$ is a constant representing static battery consumption, $K_{d}$ is a coefficient related to battery consumption due to motion, $V_{t}$ is the robot’s velocity at time step $t$, $K_{o}$ is a coefficient representing the increase in battery consumption caused by static obstacles that force the robot to modify its behaviour unexpectedly. $K_{c}$ is the charging rate when the robot is in charging mode. The two contextual factors, $C$ and $O$, indicate the robot’s charging status and the presence of a static obstacle, respectively. $C$ is 0 when the robot is discharging and $1$ when it is charging, while $O$ is 0 when no obstacle is present and $1$ when an obstacle is detected.
The $K_{s}$ and $K_{d}$ parameters were calibrated to be consistent with the official TIAGo robot’s documentation [^11] TIAGo datasheet: https://static1.squarespace.com/static/5ab0712975f9ee24a7be79df/t/5ba9cc4f652deac39f368c37/1537854552932/Datasheet_TIAGo-Hardware-Software.pdf, which states a typical battery duration of $4$ to $5$ hours. We used these manufacturer specifications to derive the parameters as follows:
  • Static consumption $K_{s}$: We assumed the $5$-hour (55 h) duration represents the robot in an idle (static) state, where sensors and systems are active but the motors are not moving. $K_{s}$ was thus calculated as the constant drain rate per second required to deplete $100\%$ of the battery in 5 hours: $K_{s}=100\%/(5~\text{h}\times 3600~\text{s/h})\approx 0.0056\%/\text{s}$
  • Dynamic Consumption $K_{d}$: We assumed the $4$-hour duration represents continuous operation at maximum velocity, $V_{\text{max}}$. We calculated $K_{d}$ by taking the total drain rate for the 4-hour duration ($100\%/(4~\text{h}\times 3600~\text{s/h})$), subtracting the static drain $K_{s}$, and then dividing by $V_{\text{max}}$: $K_{d}\approx 0.0027\%/\text{m}$.
  • Obstacle penalty $K_{o}$: This parameter is not specified in the official documentation, as it models the non-standard cost of unexpected manoeuvrers (e.g., sudden braking). This coefficient was set empirically to 3.0, representing an increase in consumption during such events.
@@
where $B_{t}$ and $B_{t-1}$ denote the battery percentage at time $t$ and t1t-1 respectively, and $L_{t-1}$ represents the battery level’s change, as modelled by (1), at time t1t-1.
@@
  • $\xleftarrow{}O\xrightarrow{}V: LL$ is affected by $O$. Moreover, $O$ influences $V$ because the robot is forced to stop or maintain a low velocity when encountering an unexpected obstacle.
@@
Although we use time-series data, we omit explicit time indices $t$ and t1t-1 for simplicity, as all the described links represent contemporaneous relationships occurring at time $t$. The considered relationships are illustrated by the Directed Acyclic Graph (DAG) in Fig. 5.
@@ 4.2.3 Worker Activities
In a warehouse, workers are assigned specific tasks at different times of the day. The activities assigned to each person are generally the same every day (Meneweger et al., 2018). For instance, warehouse operators may be responsible for unloading goods in the morning, sorting and shelving items during the day, and preparing shipments in the evening. These tasks are generally repeated every working day. As a result, certain areas of the warehouse experience higher worker traffic during specific time periods, creating congestion zones.
In order to model this phenomenon, we divided the environment into distinct areas (waypoints) and introduced time as a contextual factor to differentiate workers’ activities at different times of the day. Each waypoint represents a specific area of the warehouse that experiences varying levels of congestion depending on the time, making it an essential contextual factor in our model.
To capture the temporal dynamics of workers activity, we divided the working period into hourly time-slots from 08:00 to 18:00, along with an additional off period. Within each time-slot, worker activities are predefined and consistent, reflecting the structured nature of warehouse operations. The 08:00–09:00 slot models workers’ arrival and the beginning of the working day. To enhance realism, we model each worker arriving at a random time within this slot before starting their assigned activity. Time-slots from 09:00 to 13:00 represent the normal morning working period, during which workers are typically present in operational areas such as shelves and offices. The 13:00–14:00 slot represents the lunch break, where workers are primarily in the canteen area. Following this, time-slots from 14:00 to 18:00 model the afternoon working period, with the 17:00–18:00 slot representing the end of the working day, when workers finish their tasks and leave the warehouse after completing their standard eight-hour shift. After 18:00, we model the off period, during which no workers are present in the warehouse.
Figure 6: Two examples of people congestion areas at different time-slots. (a) Probability distribution of waypoint choices for people in time-slot 22 (09:00–10:$0$). (b) RViz view showing people primarily concentrated at the waypoints with the highest probabilities from the distribution graph. (c) Probability distribution of waypoint choices for people during lunchtime ($S6$, 13:00–14:$0$). (d) RViz view showing the majority of workers gathered in the canteen area.
An example of a congestion zone in two different time-slots is illustrated in Fig. 6. Based on this, we defined a variable called people density ($D$) to quantify the number of people in a specific waypoint. For each waypoint $W$ in the set of waypoints $\Omega$, the density $D$ is defined as follows:
@@
By incorporating both waypoints and time-slots as contextual factors, our model effectively represents fluctuating congestion levels within the warehouse. This enables a more accurate simulation of human-robot interactions, where robot navigation decisions account for dynamically changing environmental conditions. Note that while the time-slot serves as the primary contextual driver for worker activity, the model is not deterministic. Indeed, each time-slot is associated with a probability distribution over the possible waypoints, not a fixed assignment, which in turn introduces stochasticity. This allows workers to occasionally select less probable destinations, simulating natural variations in behaviour such as taking unscheduled breaks or moving between different areas, as shown in Figs. 6(a) and 6(b). In the example, you can notice that, although it is a “working time” time-slot where the working area has the highest probability of being occupied, a number of workers can still be observed in the canteen and corridors, demonstrating the non-deterministic nature of the simulation. Following this, the causal influence from the contextual factors time-slot $S$ and waypoint $W$ on the people density is $W\xrightarrow{}D\xleftarrow{}S$. Again, for simplicity, we have omitted the explicit time indices $t$ and t1t-1. The latter causal relations are added to the model in Fig. 5.
@@ 4.3 Causality-based Robot Decision Making in PeopleFlow
Our causality-based decision making framework presented in Sec. 3.3 is a general causal inference approach adaptable to various scenarios. Here, we provide more details on the factors that we estimate using causal inference and how they are used in the A* algorithm to decide whether to execute or not a given task, and in case to determine the best path.
@@
Given the robot and workers features extracted from the simulator, as well as the context factors explained in Sec. 4.2.1, the crucial aspects affecting robot task completion are the robot battery consumption $L$ and the people density $D$ in specific areas of the environment. Given their dependencies with other variables and/or contextual factors, the robot can infer $L$ based on its velocity $V$ and charging status $C$, while $D$ can be estimated using the specific time-slot $S$ and waypoint $W$. These inferences are made under the assumption that the robot has knowledge of its own velocity and charging status. Additionally, in a warehouse shared with workers, we assume that the robot is aware of the current time, enabling it to determine $S$, as well as the predefined waypoints in the environment. This allows for a probabilistic prediction of congested areas based on context (time and location), rather than a simple reactive forecast based only on past density values.
For instance, consider a scenario where the robot is assigned the task “transport an object to the delivery point,” as depicted in Fig. 2. To determine in advance whether the task can be successfully completed or if it should be immediately refused, the robot poses the following query:
“What if I go to the delivery point now at velocity vv?”
@@
To clarify, the robot queries the causal inference engine to determine whether it can start the task at the current $S$ and travel at velocity $v$. Hence, for each waypoint in $\Omega$, these two interventions allow the causal inference engine to estimate the expected values of $L$ and $D$, which represent the cost associated with the specific waypoint in terms of battery consumption and the risk of encountering people or getting stuck, respectively.
It is important to note the different nature of these two queries. The estimation of people density, $\text{P}(D=d\mid do(S=s),W=w)$, can be simplified to a standard conditional probability, $\text{P}(D=d\mid S=s,W=w)$, because there is no confounding variable between time-slot ($S$) and people density ($D$) in our model (see Fig. 5). In contrast, estimating battery consumption ($L$) requires a causal approach. The relationship between velocity ($V$) and battery level ($L$) is confounded by the presence of unexpected obstacles ($O$). Since the robot does not know in advance if it will encounter an obstacle, it cannot simply condition on $O$. Therefore, a proper causal adjustment using do-calculus is essential, motivating our choice of a causal inference engine over a standard Bayesian method.
These estimates are then used as cost components for navigation planning. In particular, the battery cost for traversing an arc between two waypoints is first computed by determining the traversal time, which is the distance between the waypoints divided by the velocity $v$ (used in the intervention). This duration is then multiplied by the expected $\hat{L}$ from Eq. (6), which corresponds to the battery consumption per time step when moving at velocity $v$. The result is the expected battery cost for that arc. These estimated values, along with the distance between waypoints, are finally used in a heuristic function defined below.
@@
$h(w_{i})=\sum\limits_{i=1}^{n-1}\Big(\lambda_{\delta}\cdot\delta(w_{i},w_{i+1})+\lambda_{D}\cdot\hat{D}(w_{i})+\lambda_{L}\cdot|\hat{L}(w_{i},w_{i+1})|\Big)$ (9)
where $\delta(w_{i},w_{i+1})$ represents the Euclidean distance between two consecutive waypoints, $\hat{D}(w_{i})$ is the expected people density at waypoint $w_{i}$ from the causal inference engine, $\hat{L}(w_{i},w_{i+1})$ is the expected battery consumption to reach waypoint $w_{i+1}$ from $w_{i}$, and $\lambda_{\delta},\lambda_{D},\lambda_{L}$ are weighting coefficients that balance the trade-off between distance minimisation, congestion avoidance, and battery efficiency.
@@
It is important to note that this decision is made once at the start of the task. We do not re-check or re-plan using the causal model during execution. This was a deliberate design choice for two reasons:
  • Crowd density is a relatively slow-moving variable. Therefore, if an area is inferred to be busy at the start of a task, it is likely to remain busy for the task’s duration. A single pre-task check is sufficient.
  • For dynamic events, such as encountering an unexpected static obstacle, we use a two-level system: (i) before the task, our battery consumption query (Eq. 4) explicitly accounts for the possibility of an obstacle. The do-calculus adjusts for the confounder $O$, so the resulting battery estimate $\hat{L}$ already considers the average cost of potential avoidance; (ii) during the task, if the robot physically encounters an obstacle, the standard ROS local planner handles the real-time avoidance.
This strategy provides a robust compromise: we make an informed, battery-aware strategic decision at the start, without incurring the high computational cost of re-running the causal query loop during execution.